OpenStudy (anonymous):

What is the remainder when P(x) = 3x^3 - x^2 + 4x - 9 is divided by x - 1?

OpenStudy (callisto):

using reminder thm, P(1) = 3(1)^3 - (1)^2 + 4(1) - 9 = -3

OpenStudy (callisto):

* reminder = P(1)